OpenStudy (needhelpinalgebra):

Ok, I got a few problems for anybody willing to help me. I will fan and medal. Explain how the Quotient of Powers was used to simplify this expression. 2^6 over 8= 2^2 By finding the quotient of the bases to be one fourth, and cancelling common factors By finding the quotient of the bases to be one fourth, and simplifying the expression By simplifying 8 to 2^3 to make both powers base two, and subtracting the exponents By simplifying 8 to 2^3 to make both powers base two, and adding the exponents

OpenStudy (needhelpinalgebra):

@Twistrose @Mr_Perfection_xD

OpenStudy (needhelpinalgebra):

@MollyNicole2015

OpenStudy (mollynicole2015):

I'm sorry, I suck at math. :\

OpenStudy (needhelpinalgebra):

Well, do you know anybody who could help me out?

OpenStudy (needhelpinalgebra):

@pink33

OpenStudy (mollynicole2015):

No, sorry. ^-^ I used to.

OpenStudy (needhelpinalgebra):

@TheSmartOne

OpenStudy (needhelpinalgebra):

@Jadeishere

OpenStudy (jadeishere):

Okay, basically.. I don't know how to explain it, actually, but i know how to solve it... so \[\frac{ 2^6 }{ 8 } = 2^2\] because a simplified version of 8 so that it comes out with the same base of 2 as the numerator is 2^3. \[\frac{ 2^6 }{ 2^3 } = 2^2\] then subtract the exponents. so it should come out to 2^3.... Um... Hm.

OpenStudy (jadeishere):

Hold on, let me go through this again and figure out what I did wrong..

OpenStudy (needhelpinalgebra):

\[2^5 \over 8 = 2^2\]

OpenStudy (needhelpinalgebra):

just not dividing the whole thing, just the first thing before the =.

OpenStudy (needhelpinalgebra):

It isn't correct, but all i need are a few quick answers XD

OpenStudy (needhelpinalgebra):

the correct answer is 4, but this isn't looking for the answer

OpenStudy (jadeishere):

Wait, oh... /).< Okay, the way it was worded confused me >.< Sorry Oh... That's so weird!! Then wouldn't it still be the third one?

OpenStudy (jadeishere):

I haven't done this math since Algebra 1 and that's been over 4 years so.. sorry

OpenStudy (needhelpinalgebra):

Its fine, i appreciate all help. I suck at this part of algebra, so... yeah.

OpenStudy (jadeishere):

Okey

OpenStudy (needhelpinalgebra):

alright. up for 2 more?

OpenStudy (jadeishere):

Sure >.<

OpenStudy (needhelpinalgebra):

actually, jk. just one.

OpenStudy (jadeishere):

Lol >.<

OpenStudy (needhelpinalgebra):

Rewrite the rational exponent as a radical.

OpenStudy (jadeishere):

Which rational exponent..

OpenStudy (needhelpinalgebra):

If you are there

OpenStudy (jadeishere):

I'm here >.< Working it out xD so what you do is set them up so it looks like you're multiplying them. \[\frac{ 3 }{ (5^4)} \times \frac{ 2 }{ 3 }\] then put them together \[\frac{ 3 * 2 }{ (5^4) * 3 }\] cancel common factors \[\frac{ 2 }{ (5^4) }\] then it'd be \[\frac{ 2 }{ 625 }\]

OpenStudy (needhelpinalgebra):

awesome. Thanks, Ma'am.

OpenStudy (needhelpinalgebra):

wait, i think you got something wrong there.

OpenStudy (jadeishere):

I did? Shoot

OpenStudy (needhelpinalgebra):

the choices i got all have to do with the square root

OpenStudy (jadeishere):

um.. 25^2? that's equal to 625

OpenStudy (jadeishere):

Eeeeeks, i dislike those..

OpenStudy (needhelpinalgebra):

I know. i do too

OpenStudy (needhelpinalgebra):

:(

OpenStudy (jadeishere):

Um... Let's see if this girl can help >.< @adrynicoleb

OpenStudy (needhelpinalgebra):

ok.

OpenStudy (jadeishere):

My oh my, i have to go! I have to go to work. Oh my god i forgot about it... i need to make up and gur... Um, tag people in it. Kay? So sorry!

OpenStudy (needhelpinalgebra):

Its fine, Ma'am. Thanks for your help. I appreciate it!!!!!

OpenStudy (needhelpinalgebra):

@ILovePuppiesLol